. Initially 100milligrams of a radioactive substance was present. After 6 hours the mass has decreased by 3%. The rate of decay is proportional to the amount of the substance present at time t. Determine the half-life of the radioactive substance.


1
Expert's answer
2021-08-24T07:54:07-0400



The rate of decay is proportional to the amount of the substance present at time t

This means that the amount of the substance can be modeled by the following differential equation:


"\frac{dQ}{dt}=-rt"


Which has the following solution:


"Q{(t)}=Q(0)e^{-rt}"


In which Q(t) is the amount after t hours, Q(0) is the initial amount and r is the decay rate.


After 6 hours the mass had decreased by 3%.

This means that . We use this to find r.


This means that "Q(6)=(1-0.03)Q(0)=0.97Q(0)"


We use this to find r


"Q{(t)}=Q(0)e^{-rt}"


"0.97Q(0)=Q(0)e^{-6r}"


"e^{-6r}=0.97"


"ln(e^{-6r})=ln(0.97)"


"-6r=ln0.97"


"r=-\frac{ln0.97}{6}"


"r=0.0051"


"Q{(t)}=Q(0)e^{-0.0051t}"



Determine the half life of radio active substance.


This is t for which "Q(t)=0.5Q(0)". So


"Q(t)=Q(0)e^{-0.0051t}"


"0.5Q(0)=Q(0)e^{-0.0051t}"


"e^{-0.0051t}=0.5"


"ln(e^{-0.0051t})=ln(0.5)"


"-0.0051t=ln0.5"


"{t}=\frac{0.5}{0.0051}"


"t=135.9"


The half-life of the radioactive substance is 135.9 hours